# Quantum Cosmology in $f(R,T)$ Theory with Schutz’s Perfect Fluid

### v1

@@ 2 $f(R,T)$ Theory and Canonical Quantization
which shows that fluid energy-momentum is not covariantly conserved in $f(R,T)$ gravity due to $f_{T}$-dependent source terms. Note that using $\mathcal{L}_{\mathrm{matt}}=-\rho$ modifies $\Theta_{\mu\nu}$ and the non-conservation law affects the minisuperspace coupling to $T$. Since Schutz’s Hamiltonian formulation is naturally compatible with $\mathcal{L}_{\mathrm{matt}}=p$, we maintain this choice throughout the paper.
@@ 4.3 The case of $G^{0}(T)=T$
$\begin{split}&\dot{a}=\{a,H\}=-N\frac{7}{5}\Big(\frac{1}{b_{2}}\Big)\frac{Pa}{24a},\\ &\dot{P}_{a}=\{P_{a},H\}=-N\Bigg[\frac{7}{5}\Big(\frac{1}{b_{2}}\Big)\frac{P_{a}^{2}}{24a^{2}}-\frac{9}{7}a^{2}b_{1}-\frac{30}{7}kb_{2}-3\alpha\frac{P_{\tau}}{a^{3\alpha+1}}\Bigg],\\ &\dot{\tau}=\{\tau,H\}=N\frac{15}{7}\frac{1}{a^{3\alpha}},\\ &\dot{P}_{\tau}=\{P_{\tau},H\}=0.\end{split}$ (88)
@@
$2b_{2}\dot{a}^{2}a^{1-3\alpha}-\frac{1}{5}b_{1}a^{3(\alpha+1)}-kb_{2}a^{1+3\alpha}+P_{\tau}=0.$ (121)
@@
$\Bigg(\frac{\dot{a}}{a}\Bigg)^{2}=\frac{1}{2b_{2}}\Bigg[\frac{1}{5}b_{1}a^{6}+kb_{2}a^{4}-P_{\tau}\Bigg].$ (123)
@@
    $\Bigg(\frac{\dot{a}}{a}\Bigg)^{2}=\frac{k}{2}\,a^{4}-\frac{P_{\tau}}{2b_{2}}.$ (124)
@@
    $a(t)=\begin{cases}\left(\sqrt{\dfrac{P_{\tau}}{b_{2}}}\;\sec\!\Big(\sqrt{\dfrac{2P_{\tau}}{b_{2}}}\,(t-t_{0})\Big)\right)^{\!1/2}&k=+1,\;\dfrac{P_{\tau}}{b_{2}}>0\\[10.0pt] \left(\dfrac{1}{\sqrt{2}\,(t-t_{0})}\right)^{\!1/2}&k=+1,\;P_{\tau}=0\\[10.0pt] \left(\sqrt{-\dfrac{P_{\tau}}{b_{2}}}\;\operatorname{cosech}\!\Big(\sqrt{-\dfrac{2P_{\tau}}{b_{2}}}\,(t-t_{0})\Big)\right)^{\!1/2}&k=+1,\;\dfrac{P_{\tau}}{b_{2}}<0\\[12.0pt] \left(\sqrt{-\dfrac{P_{\tau}}{b_{2}}}\;\operatorname{sech}\!\Big(\sqrt{-\dfrac{2P_{\tau}}{b_{2}}}\,(t-t_{0})\Big)\right)^{\!1/2}&k=-1,\;\dfrac{P_{\tau}}{b_{2}}\leq 0\\[10.0pt] \text{no real solution}&k=-1,\;\dfrac{P_{\tau}}{b_{2}}>0.\end{cases}$ (125)
@@
    $\Bigg(\frac{\dot{a}}{a}\Bigg)^{2}=\frac{1}{2b_{2}}\Bigg[\frac{1}{5}b_{1}a^{6}-P_{\tau}\Bigg].$ (126)
@@
    $a(t)=\begin{cases}\left(\sqrt{\dfrac{5P_{\tau}}{b_{1}}}\,\sec\!\Big(3\sqrt{\dfrac{{P_{\tau}}}{{2b_{2}}}}\,(t-t_{0})\Big)\right)^{\!1/3}&\;P_{\tau}>0\\[10.0pt] \left(\pm 3\sqrt{\dfrac{b_{1}}{10b_{2}}}\,(t-t_{0})\right)^{-1/3}&\;P_{\tau}=0\\[10.0pt] \left(\sqrt{-\dfrac{5P_{\tau}}{b_{1}}}\,\cosh\!\Big(3\sqrt{\dfrac{{-P_{\tau}}}{{2b_{2}}}}\,(t-t_{0})\Big)\right)^{\!1/3}&\;P_{\tau}<0.\end{cases}$ (127)
    Here, $t_{0}$ is an integration constant determined by the initial conditions. Moreover, real solutions require the right-hand side of (126) to be non-negative.
@@
    $\Bigg(\frac{\dot{a}}{a}\Bigg)^{2}=-\frac{P_{\tau}}{2b_{2}}.$ (128)
@@
[…]
@@
$=$ $\frac{1}{(2\pi\sigma^{2})^{1/4}}\int_{-\infty}^{\infty}e^{-\frac{(k^{\prime}-k_{0})^{2}}{4\sigma^{2}}+i(k^{\prime}x-\mathcal{E}\tau)}dk^{\prime}$
$=$ $\frac{1}{(2\pi\sigma^{2})^{1/4}}\int_{-\infty}^{\infty}e^{-(\frac{1}{4\sigma^{2}}+\frac{it}{2m})k^{\prime 2}+(\frac{k_{0}}{2\sigma^{2}}+ix)k^{\prime}+(i\vartheta\tau-\frac{k_{0}^{2}}{4\sigma^{2}})}dk^{\prime}$
$=$ $\frac{1}{(2\pi\sigma^{2})^{1/4}}\sqrt{\frac{\pi}{\frac{1}{4\sigma^{2}}+\frac{it}{2m}}}\>e^{\frac{(\frac{k_{0}}{2\sigma^{2}}+ix)^{2}}{4(\frac{1}{4\sigma^{2}}+\frac{it}{2m})}+(i\vartheta\tau-\frac{k_{0}^{2}}{4\sigma^{2}})}.$
@@
$\varPhi(x,\tau)\propto e^{\frac{(\frac{k_{0}}{2\sigma^{2}}+ix)^{2}}{4(\frac{1}{4\sigma^{2}}+\frac{it}{2m})}},$ (136)
@@
$\langle a(\tau)\rangle=\mathbb{E}[e^{x}]=e^{\mu+\sigma_{x}/2}=\exp\!\left(-\frac{2k_{0}}{c_{2}}\,\tau+\frac{1}{8\sigma^{2}}+\frac{2\sigma^{2}\tau^{2}}{c_{2}^{2}}\right).$ (137)
@@
$\ln\langle a(\tau)\rangle\sim\frac{2\sigma^{2}}{c_{2}^{2}}\,\tau^{2},\qquad\langle a(\tau)\rangle\propto\exp\!\left(\frac{2\sigma^{2}}{c_{2}^{2}}\tau^{2}\right),$ (140)

### v2

@@ 2 $f(R,T)$ Theory and Canonical Quantization
which shows that fluid energy-momentum is not covariantly conserved in $f(R,T)$ gravity due to $f_{T}$-dependent source terms. Note that using $\mathcal{L}_{\mathrm{matt}}=-\rho$ modifies $\Theta_{\mu\nu}$ and the non-conservation law affects the minisuperspace coupling to $T$. Since Schutz’s Hamiltonian formulation is naturally compatible with $\mathcal{L}_{\mathrm{matt}}=p$, we maintain this choice throughout the paper. See also [24], where the gravitational field equations for theories with trace couplings to a relativistic perfect fluid were rederived using both the Brown and Schutz variational formalisms, and it was noted that certain results in the literature may benefit from further careful analysis.
@@
$\dot{D}$ $=$ $\{D,H\}=-\frac{N}{6a^{3}}\left(\frac{\mathcal{A}}{B}P_{M}\right),$
$\dot{E}$ $=$ $\{E,H\}=-\frac{N}{6a^{3}}\left(\frac{\mathcal{A}}{B}\frac{1}{3}P_{F}E\right),$
@@ 4.3 The case of $G^{0}(T)=T$
$\begin{split}&\dot{a}=\{a,H\}=-N\frac{7}{5}\Big(\frac{1}{b_{2}}\Big)\frac{P_{a}}{12a},\\ &\dot{P}_{a}=\{P_{a},H\}=-N\Bigg[\frac{7}{5}\Big(\frac{1}{b_{2}}\Big)\frac{P_{a}^{2}}{24a^{2}}-\frac{9}{7}a^{2}b_{1}-\frac{30}{7}kb_{2}-3\alpha\frac{P_{\tau}}{a^{3\alpha+1}}\Bigg],\\ &\dot{\tau}=\{\tau,H\}=N\frac{15}{7}\frac{1}{a^{3\alpha}},\\ &\dot{P}_{\tau}=\{P_{\tau},H\}=0.\end{split}$ (88)
@@
$-2b_{2}\dot{a}^{2}a^{1-3\alpha}-\frac{1}{5}b_{1}a^{3(\alpha+1)}-2kb_{2}a^{1+3\alpha}+P_{\tau}=0.$ (121)
@@
$\Bigg(\frac{\dot{a}}{a}\Bigg)^{2}=\frac{-1}{2b_{2}}\Bigg[\frac{1}{5}b_{1}a^{6}+2kb_{2}a^{4}-P_{\tau}\Bigg].$ (123)
@@
    $\Bigg(\frac{\dot{a}}{a}\Bigg)^{2}=-{k}\,a^{4}+\frac{P_{\tau}}{2b_{2}}.$ (124)
@@
    $a(t)=\begin{cases}\left(\sqrt{\dfrac{P_{\tau}}{2b_{2}}}\;\operatorname{sech}\!\Big(\sqrt{\dfrac{2P_{\tau}}{b_{2}}}\,(t-t_{0})\Big)\right)^{\!1/2}&k=+1,\;\dfrac{P_{\tau}}{b_{2}}>0\\[12.0pt] \text{no real solution}&k=+1,\;\dfrac{P_{\tau}}{b_{2}}\leq 0\\[12.0pt] \left(\sqrt{\dfrac{P_{\tau}}{2b_{2}}}\;\operatorname{cosech}\!\Big(\sqrt{\dfrac{2P_{\tau}}{b_{2}}}\,(t-t_{0})\Big)\right)^{\!1/2}&k=-1,\;\dfrac{P_{\tau}}{b_{2}}>0\\[12.0pt] \left(\dfrac{1}{2\,(t-t_{0})}\right)^{\!1/2}&k=-1,\;P_{\tau}=0\\[12.0pt] \left(\sqrt{-\dfrac{P_{\tau}}{2b_{2}}}\;\sec\!\Big(\sqrt{-\dfrac{2P_{\tau}}{b_{2}}}\,(t-t_{0})\Big)\right)^{\!1/2}&k=-1,\;\dfrac{P_{\tau}}{b_{2}}<0.\end{cases}$ (125)
@@
    $\Bigg(\frac{\dot{a}}{a}\Bigg)^{2}=\frac{-1}{2b_{2}}\Bigg[\frac{1}{5}b_{1}a^{6}-P_{\tau}\Bigg].$ (126)
@@
    $a(t)=\begin{cases}\left(\sqrt{\dfrac{5P_{\tau}}{b_{1}}}\,\operatorname{sech}\!\Big(3\sqrt{\dfrac{{P_{\tau}}}{{2b_{2}}}}\,(t-t_{0})\Big)\right)^{\!1/3}&\;P_{\tau}>0\\[12.0pt] \text{no real solution}&\;P_{\tau}\leq 0.\end{cases}$ (127)
@@
    $\Bigg(\frac{\dot{a}}{a}\Bigg)^{2}=\frac{P_{\tau}}{2b_{2}}.$ (128)
@@
    $a(t)=\begin{cases}a_{0}\,\exp\!\left[\pm\sqrt{\dfrac{P_{\tau}}{2b_{2}}}\;(t-t_{0})\right]&\dfrac{P_{\tau}}{b_{2}}>0\\[6.0pt] a_{0}&{P_{\tau}}=0\\[6.0pt] \text{no real solution}&\dfrac{P_{\tau}}{b_{2}}<0.\end{cases}$ (129)
@@
$=$ $\frac{1}{(2\pi\sigma^{2})^{1/4}}\int_{-\infty}^{\infty}\exp\!\left[-\frac{(k^{\prime}-k_{0})^{2}}{4\sigma^{2}}+i(k^{\prime}x-\mathcal{E}\tau)\right]\,dk^{\prime}$
$=$ $\frac{1}{(2\pi\sigma^{2})^{1/4}}\int_{-\infty}^{\infty}\exp\!\left[-\left(\frac{1}{4\sigma^{2}}+\frac{i\tau}{2m}\right)k^{\prime 2}+\left(\frac{k_{0}}{2\sigma^{2}}+ix\right)k^{\prime}+\left(-i\vartheta\tau-\frac{k_{0}^{2}}{4\sigma^{2}}\right)\right]\,dk^{\prime}$
$=$ $\frac{1}{(2\pi\sigma^{2})^{1/4}}\sqrt{\frac{\pi}{\frac{1}{4\sigma^{2}}+\frac{i\tau}{2m}}}\>\exp\!\left[\frac{\left(\frac{k_{0}}{2\sigma^{2}}+ix\right)^{2}}{4\left(\frac{1}{4\sigma^{2}}+\frac{i\tau}{2m}\right)}+\left(-i\vartheta\tau-\frac{k_{0}^{2}}{4\sigma^{2}}\right)\right].$
@@
$\varPhi(x,\tau)\propto\exp\!\left[\frac{\left(\frac{k_{0}}{2\sigma^{2}}+ix\right)^{2}}{4\left(\frac{1}{4\sigma^{2}}+\frac{i\tau}{2m}\right)}\right].$ (136)
@@
$\langle a(\tau)\rangle=\mathbb{E}[e^{x}]=e^{\mu+\sigma_{x}^{2}/2}=\exp\!\left(-\frac{2k_{0}}{c_{2}}\,\tau+\frac{1}{8\sigma^{2}}+\frac{2\sigma^{2}\tau^{2}}{c_{2}^{2}}\right).$ (137)
@@
$\ln\langle a(\tau)\rangle\sim\frac{2\sigma^{2}}{c_{2}^{2}}\,\tau^{2},\qquad\langle a(\tau)\rangle\propto\exp\!\left(\frac{2\sigma^{2}}{c_{2}^{2}}\tau^{2}\right).$ (140)